OpenStudy (anonymous):

Find the distance between the pair of points (-18, -9) and (-4, 21).

OpenStudy (anonymous):

a. 25.06 b. 36.8 c. 33.11 d. 26.57

OpenStudy (anonymous):

\[\sqrt{(-18-(-4))^2 + (-9-21)^2}\] \[\sqrt{(-14)^2 + (-30)^2}\] \[\sqrt{196 + 900}\] = 33.11